OpenStudy (anonymous):

given sin theta= -3/5 and csc theta= -5/3 in quadrant III, find the value of cos theta and sec theta

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

\[\sin(\theta)=-\frac{3}{5}\] means \[\csc(\theta)=-\frac{5}{3}\] this is one piece of information not two

OpenStudy (anonymous):

you have the world famous 3 - 4 - 5 right triangle in quadrant 3 

OpenStudy (anonymous):

you can take any trig ratio you like

OpenStudy (anonymous):

@satellite73 so cos= -4/5 and sec= -5/4?